which if the followings has the shortest length
400pm
4.0 x10^-4m
4.0 x 10^-10cm
4.0 nm

how many atoms are in 236 grams of arsenic?
0.0695
9.51 x 10^22
4.19 x 10^22
1.90 x 10^24
7.80 x 10^24


which of the following is iconic compound?
kcl
co2
no2
pcl3
Answer #1

1)

(a) 1pm = 10-12m.

Therfore, 400 pm = 4 \times 10-10 m.

(b) 4.0 \times 10-4m

(c) 4.0 \times 10-10cm = 4.0 \times 10-12cm.

(d) 1nm = 10-9m.

Therfore, 4.0 nm= 4.0\times​​​​​​10-9m.

Therefore, option (c) is the shortest of all.

2)

Arsenic atomic weight = 74.92 g/mol.

Therefore, no.of moles of arsenic present= 236 g / 74.92 g/mol = 3.15 mol.

And 1 mole of aresnic contains 6.023\times 1023 atoms of arsenic.

Therefore, total no.of arsenic atoms present in 236 g of arsenic = 3.15\times 6.023\times 1023=1.90 \times 1024.

3)

An ionic compound is the compound in which the atoms are ions and are held by the electrostatic force of attraction.Like there are positive and negative charges are held by balanced electrostatic forces.

Whereas in covalent compounds, the atoms are hepd together by covalent bonds like the sigma bond, pi bond which we study, where the orbitals overlap 2ith each other.

In the given options, only KCl is the ionic compound and the others are covalent.